OpenStudy (anonymous):

Find the 23rd term of the arithmetic sequence with the terms a1=27 and d=16

OpenStudy (anonymous):

t23=a1+22d

OpenStudy (anonymous):

t23=27+22*16=379

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

U r welcome!